OpenStudy (anonymous):

2w-y=7w-s, solve for w I really dont get this, someone please help

OpenStudy (anonymous):

For the s its actually 2

OpenStudy (anonymous):

So switch out the s for 2

OpenStudy (anonymous):

2w-y=7w-2 +2 +2

OpenStudy (anonymous):

so 2-y/5=5/5w

OpenStudy (anonymous):

w=(2-y)/5

OpenStudy (anonymous):

i hope this helps

OpenStudy (anonymous):

it does thank you

OpenStudy (anonymous):

welcome